OpenStudy (anonymous):

Find the 10th term in the Arithmetic Sequence -2, -5, -8, -11 A) -10 B) -29 C) 22 D) -21 @malcolmmcswain

OpenStudy (malcolmmcswain):

Ok, what seems to be the pattern?

OpenStudy (anonymous):

i can help :)

OpenStudy (malcolmmcswain):

Are we increasing/decreasing? What is our sequence?

OpenStudy (anonymous):

Increasing

OpenStudy (malcolmmcswain):

No, we are moving backwards in the negatives.

OpenStudy (anonymous):

wait i'm confused now

OpenStudy (malcolmmcswain):

And by a specific number.

OpenStudy (anonymous):

when u keep going on with the pattern it goes to -30

OpenStudy (malcolmmcswain):

We are moving backwards by one number I.E. n - a number

OpenStudy (malcolmmcswain):

What number is that? Hint -5 - a number = -8

OpenStudy (tkhunny):

1st term: -2 2nd term: -5 = -2 - 3 3rd term: -8 = -2 - 3(2) 4th term: -11 = -2 - 3(3) ... 10th term: ?? = -2 - 3(9) THAT'S the pattern.

OpenStudy (anonymous):

would it be B?

OpenStudy (malcolmmcswain):

Congrats on 1000 fans, tk. :)

OpenStudy (malcolmmcswain):

-29 = -2 - 3(9) Correct. :D

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

B